# Why Smishing Works

### v2

@@ 6.5. Limitations
Our study has several limitations that warrant acknowledgment. Firstly, we note the disproportionate number of iOS screenshots compared to Android, with only two out of 16 screenshots being Android-based. This potential bias toward iOS users could impact the generalizability of our results, considering the influence of sampling bias on survey validity, as highlighted by Royal et al. (royal2019survey). However, we cannot definitively conclude whether participants’ preference for a particular mobile OS significantly influenced their ability to detect smishing SMSes accurately in our survey.
Secondly, we acknowledge the potential compromise in results due to simulated study conditions (schechter2007emperor; lain2022phishing). Participants viewed screenshots within an online Qualtrics survey rather than messages directly from their messaging app. However, this approach was chosen to prioritize participant security and privacy. Additionally, it aimed to eliminate bias stemming from differences in mobile applications, SMS filtering systems, and security indicators on various devices. Thus, each user was provided with the same viewing context for each message. Lastly, it is important to recognize that in a setting where participants are aware of a security study being conducted, they tend to be more cautious. To mitigate this effect, we initially inquired about participants’ likelihood to interact with and their comfort level in interacting with all messages before introducing the topic of real or fake messages at a later stage of the study.
Lastly, we recognize the potential impact of survey fatigue, given the extensive nature of our survey. It comprised a pre-survey demographics questionnaire, 32 multi-part main survey questions, and five post-survey questionnaires. Fatigue may have affected participants’ attention, potentially influencing result accuracy. Nevertheless, the substantial number of responses received suggests that participants engaged thoughtfully with our survey, and a high level of interaction can contribute to result validity (wei2023skilled).
@@ 6.6. Future Work
In a subsequent study, we aim to apply the principles established in this research, tailoring them specifically to vulnerable populations such as the elderly. Unfortunately, the elderly are more susceptible to online fraud compared to younger demographics (FBI2022). Individuals over the age of 60 experienced higher financial losses to fraud in 2022, losing a total of \$3.1 billion, which is 1.7 times more than the combined losses of individuals under the age of 20, 20-29, and 30-39, according to the 2022 FBI Elder Fraud report (FBI2022). Given these alarming losses, it is essential further to investigate the perceptions of the elderly towards smishing. This effort aims to develop effective methods to protect them from the increasing threat of smishing attacks.
@@ 8. Conclusion
We conducted a smishing detection study involving 214 participants to unravel the factors influencing users’ accuracy with both real and fake messages. Our findings indicate a smishing detection rate of 65.6% for fake messages, contrasting with a lower 44.6% for real SMSes. Additionally, we pinpointed areas of interest in messages that significantly affected users’ accuracy in identifying both real and fake messages. These results underscore the considerable room for improvement in how brands communicate messages to their customers. Our findings will contribute to the development of warning systems for smishing and the creation of more effective educational tools aimed at increasing user awareness of identifying real and fake messages.
9. Acknowledgement
ChatGPT was employed to generate sections of this work, such as graphs and code. It also served to correct grammatical mistakes and improve the overall clarity and readability of the text. For this purpose, we used the following prompt: ”Please rectify grammatical errors and enhance the readability and clarity of this paragraph.”
Appendix

### v4

@@ 6.5. Limitations
Our study has several limitations that warrant acknowledgment. Firstly, we note the disproportionate number of iOS screenshots compared to Android, with only two out of 16 screenshots being Android-based. This potential bias toward iOS users could impact the generalizability of our results, considering the influence of sampling bias on survey validity, as highlighted by Royal et al. (royal2019survey, ). However, we cannot definitively conclude whether participants’ preference for a particular mobile OS significantly influenced their ability to detect smishing SMSes accurately in our survey.
Secondly, we acknowledge the potential compromise in results due to simulated study conditions (schechter2007emperor, ; lain2022phishing, ). Participants viewed screenshots within the Qualtrics survey, rather than directly from their messaging app. This approach was chosen to prioritize participant security and privacy. Additionally, it aimed to eliminate bias by providing with the same viewing context for each message. Lastly, it is important to recognize that in a setting where participants are aware of a security study being conducted, they tend to be more cautious. To mitigate this effect, we initially inquired about participants’ likelihood to interact with and their comfort level in interacting with all messages before introducing the topic of real or fake messages at a later stage of the study.
Additionally, to measure to elements of the message content, we intentionally excluded the context of whether a message was expected or triggered by user activity. While this information can be important for distinguishing real from fake messages, we believe the necessary information to determine a message’s authenticity can be found within the content itself. This approach allowed us to focus attention on the AOI in the message, acknowledging that this limitation may affect the realism of the study.
Lastly, we recognize the potential impact of survey fatigue, given the extensive nature of our survey. It comprised a pre-survey demographics questionnaire, 32 multi-part main survey questions, and five post-survey questionnaires. Fatigue may have affected participants’ attention, potentially influencing result accuracy. Nevertheless, the substantial number of responses received suggests that participants engaged thoughtfully with our survey, and a high level of interaction can contribute to result validity (wei2023skilled, ).
@@ 6.6. Future Work
In a subsequent study, we aim to apply the principles established in this research to vulnerable populations, such as the elderly. Unfortunately, the elderly are more susceptible to online fraud compared to younger demographics (FBI2022, ). Individuals over the age of 60 experienced higher financial losses to fraud in 2022, losing a total of \$3.1 billion, which is 1.7 times more than the combined losses of individuals under the age of 20, 20-29, and 30-39, according to the 2022 FBI Elder Fraud report (FBI2022, ). Given these alarming losses, it is essential further to investigate the perceptions of the elderly towards smishing. This effort can help develop effective methods to protect them from the increasing threat of smishing attacks. Finally, based on our results we found significant differences between the likelihood and comfort in interacting with our smishing messages in various populations. Future work could explore potential socioeconomic factors that may contribute to this increase.
@@ 8. Conclusion
We conducted a smishing detection study involving 187 participants to unravel the factors influencing users’ accuracy with both real and fake messages. Our findings indicate a smishing detection rate of 67.1% for fake messages, contrasting with a lower 43.6% for real SMSes. Additionally, we pinpointed areas of interest in messages that significantly affected users’ accuracy in identifying both real and fake messages. These results underscore the considerable room for improvement in how brands communicate messages to their customers. Our findings will contribute to the development of warning systems for smishing and the creation of more effective educational tools aimed at increasing user awareness of identifying real and fake messages.